A pile driver hammer of mass 140 kg falls freely through a distance of 4.5 m to strike a pile of mass 390 kg and drives it 70 mm into the ground. The hammer does not rebound when driving the pile. Determine the average resistance of the ground.


1
Expert's answer
2021-02-24T07:05:52-0500

The potential energy of the pile-driver is converted into kinetic energy. Thus


"\text{potential energy} = \text{kinetic energy,}""mgH=\dfrac{mv^2}{2}"

Find velocity of pile immediately after impact using principle of conservation of momentum


"mv=(M+m)u"

"u=\dfrac{m}{M+m}v"

The pile-driver and pile together have a mass "(m+M)" and possess kinetic energy "\dfrac{(M+m)u^2}{2}."

The change in potential energy of the driver as it moves in the pile and the pile as it moves through the ground


"(M+m)g(0-h)"


Let "R=" the average resistance of the ground. Then


"\text{work done}=R\times h"

By the principle of conservation of energy


"\dfrac{(M+m)u^2}{2}=(M+m)g(0-h)+R\times h"

"R=(M+m)g+\dfrac{(M+m)u^2}{2h}"

"R=(M+m)g+\dfrac{m^2v^2}{2h(M+m)}"

"R=(M+m)g+\dfrac{m^2gH}{h(M+m)}"

"(140kg+390kg)(9.81m/s^2)+\dfrac{(140kg)^2(9.81m/s^2)(4.5m)}{0.07m(140kg+390kg)}"

"R=28521N"


The average resistance of the ground is 28.521 kN.

We see that the kinetic energy of the system at impact is not conserved.